OpenStudy (anonymous):

If the electric field has magnitude P 3.00m from a +q charge what is the magnitude of the electric field 9.00m from the charge? .(+q)______x(3.00m)_________x(9.00m)

OpenStudy (anonymous):



OpenStudy (anonymous):

E = k q / r^2 E1/E2 = (r2/r1)^2

OpenStudy (anonymous):

so my electric field magnitude at 9.00m from charge is : E=kq/9^2=kq/81??